# Sharp Analysis for KL-Regularized Contextual Bandits and RLHF

### v1

@@ 2.1 Contextual Bandits with KL Regularization
$Q(\pi)=\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi(\cdot|x)}\bigg{[}R(\theta_{*},x,a)-\eta^{-1}\ln\frac{\pi(a|x)}{\pi_{0}(a|x)}\bigg{]},$ (2.1)
@@
2.2 Reinforcement Learning from Preference Feedback
The second framework we consider is preference feedback, which is widely applied in projects such as Chat-GPT (OpenAI, 2023) and Claude (Bai et al., 2022). Specifically, when receiving a prompt $x\in\mathcal{X}$, and two actions (responses) $a^{1},a^{2}\in\mathcal{A}$ from some LLM policy $\pi(\cdot|x)$, a preference oracle will give feedback $y$ defined as follows:
Definition 2.2 (Preference Oracle).
Definition 2.3 (Bradley-Terry Model).
where $\sigma(\cdot)$ is the sigmoid function.
2.3 Additional Notations and Definitions
In this subsection, we introduce the definitions shared by both settings.
Reward function class.
@@ 2.3 Additional Notations and Definitions
Hence, the comparator policy is the solution to the oracle given the true reward function $R(\theta^{*},\cdot,\cdot)$: $\pi^{*}(\cdot|\cdot)\propto\pi_{0}(\cdot|\cdot)\cdot\exp(\eta R(\theta^{*},\cdot,\cdot))$. The goal is to minimize the sub-optimality of our learned policy $\widehat{\pi}$ with respect to $\pi^{*}$: $Q(\pi^{*})-Q(\widehat{\pi})$.
Coverage conditions.
It is crucial to assume that our data-collector policy $\pi_{0}$ possesses good coverage, which can ensure that the learned reward function can generalize well to unseen contexts (prompts) and actions (responses), and thus can enable us to approximate the optimal policy. The global coverage is the uniform cover over all the policies in the considered class $\Pi$, which is standard in offline RL (Munos and Szepesvári, 2008; Song et al., 2024) and online RL (Xie et al., 2022; Rosset et al., 2024). Essentially, Song et al. (2024) demonstrated that global coverage is necessary for offline framework and Direct Preference Optimization (DPO) fails without global coverage. Hence, we introduce two types of global-coverage conditions.
@@
Given a reference policy $\pi_{0}$, $D^{2}$ is the minimum positive real number satisfying $\forall\ {(x,a)\in\mathcal{X}\times\mathcal{A},s.t.\ \pi(a|x)>0}$ and $\forall\ b:\mathcal{X}\to[-B,B]$, we have
$\sup_{\theta,\theta^{\prime}\in\Theta}\frac{|R(\theta^{\prime},x,a)-R(\theta,x,a)-b(x)|^{2}}{\mathbb{E}_{x^{\prime}\sim d_{0}}\mathbb{E}_{a^{\prime}\sim\pi_{0}(\cdot|x^{\prime})}|R(\theta^{\prime},x^{\prime},a^{\prime})-R(\theta,x^{\prime},a^{\prime})-b(x^{\prime})|^{2}}\leq D^{2}.$
The coverage coefficient $D$ measures how well the in-sample error induced by distribution $d_{0}\times\pi_{0}$ can cover the out-of-sample error, i.e., it depicts the ability of $\pi_{0}$ to cover the action space. This concept is adapted from the F-design for online RL under general function approximation (Agarwal et al., 2024), and resembles the coverage coefficient for offline RL (Ye et al., 2024c, a), and the eluder dimension (Wang et al., 2020; Ye et al., 2023; Agarwal et al., 2023; Zhao et al., 2023a) for online RL.
Definition 2.7 (Global-Policy Coverage).
The two conditions require the reference policy to cover all possible policy distributions, which is standard and common in RL literature. Additionally, although the two conditions defined above are both global, it is obvious that $D^{2}\leq C_{\mathrm{GL}}$, indicating that it is more general to assume a finite $D$.
Because of the KL-regularization for RLHF, the learned policy will not move too far from the reference policy. Hence, it is natural to relax the global coverage to local coverage inside the KL-ball (Song et al., 2024).
Definition 2.8 (Local KL-ball Coverage).
Remark 2.9 (Relation between Local and Global Coverage Conditions).
3 KL-Regularized Contextual Bandits
@@ 3.1 Lower Bound
In this section, we provide a lower bound for the KL-regularized contextual bandit problem.
@@
For any $\epsilon\in(0,1),\eta>0$, and any algorithm $A$, there exists a KL-regularized contextual bandit problem with $O(1)$ coverage coefficient and reward function class $\mathcal{R}$ such that $A$ requires at least $\Omega\big{(}\min(\frac{\eta\log N_{\mathcal{R}}(\epsilon)}{\epsilon},\frac{\log N_{\mathcal{R}}(\epsilon)}{\epsilon^{2}})\big{)}$ rounds to achieve a suboptimality gap of $\epsilon$.
@@
The lower bound in Theorem 3.1 indicates that the sample complexity of the KL-regularized contextual bandit problem is $\Omega(\eta\log N_{\mathcal{R}}(\epsilon)/\epsilon)$ when $\epsilon$ is sufficiently small. In our proof, the KL-regularization term shifts the local landscape of the objective function, which prevents us to directly apply the standard bandit analysis, and thus requires a novel analysis to derive the new lower bound. This $\Omega(\eta\log N_{\mathcal{R}}(\epsilon)/\epsilon)$ lower bound suggests that the KL-regularized contextual bandit problem enjoys a lower sample complexity compared to the standard contextual bandit.
@@ 3.2 The Proposed Algorithm
Algorithm 1 Two-Stage mixed-policy sampling
@@
10:for $i$ = $1,\ldots,n$ do
11:     Sample context $x_{i}\sim d_{0}$ and action $a_{i}\sim\pi_{\widehat{\theta}_{0}}^{\eta}(\cdot|x_{i})$.
12:     Observe reward $r_{i}=R(\theta_{*},x_{i},a_{i})+\epsilon_{i}$, where $\epsilon_{i}$ is the random noise.
13:end for
14:Compute the least square estimate of the reward function using $\{(x_{i},a_{i},r_{i})\}_{i=1}^{n}$ together with $D_{0}$:
@@
We present the algorithmic framework in Algorithm 1 for the KL-regularized contextual bandit problem, which serves as a theoretical model for online RLHF with absolute-rating feedback. The algorithm consists of two states:
  • In the second stage, we utilize the trained policy $\pi_{\widehat{\theta}_{0}}^{\eta}$ to sample $n$ contexts (prompts) and actions (responses). With the intermediate policy $\pi_{\widehat{\theta}_{0}}^{\eta}$, we can collect new data $\{(x_{i},a_{i},r_{i})\}_{i=1}^{n}$ which is more aligned with the data distribution induced by the optimal policy $\pi_{*}$. In line 14, the algorithm combines data from both stages $\{(x_{i},a_{i},r_{i})\}_{i=1}^{n}$ and $\{(x_{i}^{0},a_{i}^{0},r_{i}^{0})\}_{i=1}^{m}$ to compute a refined least squares estimate $\widehat{\theta}$ of the reward function, minimizing the sum of squared errors across both datasets. By aggregating the two datasets together, there is an overlap between the data to compute $\widehat{\theta}$ and $\widehat{\theta}_{0}$, so that the output policy $\pi_{\widehat{\theta}}^{\eta}$ is well covered by the intermediate policy $\pi_{\widehat{\theta}_{0}}^{\eta}$.
@@ 3.3 Theoretical Guarantees
Revisiting Previous Analysis.
The previous analysis (e.g. Xiong et al., 2024a) is not tight since they basically follow the techniques of bandits and neglect the significance of KL-regularization. For simplicity, We use short-hand notation $R(\theta,x,\pi)=\mathbb{E}_{a\sim\pi(\cdot|x)}R(\theta,x,a)$ and denote $\mathrm{KL}(\pi(\cdot|x)\|\pi^{\prime}(\cdot|x))$ by $\mathrm{KL}(\pi\|\pi^{\prime})$ when there is no confusion. We make the estimation on a dataset $\{(x_{i},a_{i},r_{i}):x_{i}\sim d_{0},s_{i}\sim\pi_{0}\}_{i=1}^{n}$: $\pi_{\widehat{\theta}}^{\eta}=\mathop{\mathrm{argmax}}_{\pi\in\Pi}\mathbb{E}_{x\sim d_{0}}[R(\widehat{\theta},x,\pi)-\eta^{-1}\mathrm{KL}(\pi\|\pi_{0})]$, and has a small in-sample-error: $\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{o}(\cdot|x)}\big{[}(R(\widehat{\theta},x,a)-R(\theta^{*},x,a))^{2}\big{]}=O(1/n).$ The sub-optimality is decomposed as:
$Q(\pi^{*})-Q(\pi_{\widehat{\theta}}^{\eta})=$ $\mathbb{E}_{x\sim d_{0}}\big{[}R(\theta^{*},x,\pi^{*})-R(\widehat{\theta},x,\pi^{*})\big{]}+\mathbb{E}_{x\sim d_{0}}\big{[}R(\widehat{\theta},x,\pi_{\widehat{\theta}}^{\eta})-R(\theta^{*},x,\pi_{\widehat{\theta}}^{\eta})\big{]}$
$+\mathbb{E}_{x\sim d_{0}}\big{[}R(\widehat{\theta},x,\pi^{*})-\eta^{-1}\mathrm{KL}(\pi^{*}\|\pi_{0})\big{]}-\mathbb{E}_{x\sim d_{0}}\big{[}R(\widehat{\theta},x,\pi_{\widehat{\theta}}^{\eta})-\eta^{-1}\mathrm{KL}(\pi_{\widehat{\theta}}^{\eta}\|\pi_{0}))\big{]}$
$\leq$ $\mathbb{E}_{x\sim d_{0}}\big{[}R(\theta^{*},x,\pi^{*})-R(\widehat{\theta},x,\pi^{*})+R(\widehat{\theta},x,\pi_{\widehat{\theta}}^{\eta})-R(\theta^{*},x,\pi_{\widehat{\theta}}^{\eta})\big{]},$
@@
$Q(\pi^{*})-Q(\pi_{\widehat{\theta}}^{\eta})\leq$ $2C_{\mathrm{GL}}\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{0}(\cdot|x)}\big{[}|R(\theta^{*},x,a)-R(\widehat{\theta},x,a)|\big{]}$
$\leq$ $2C_{\mathrm{GL}}\sqrt{\mathbb{E}_{a\sim\pi_{0}(\cdot|x)}\big{[}(R(\theta^{*},x,a)-R(\widehat{\theta},x,a))^{2}\big{]}}=O(C_{\mathrm{GL}}/\sqrt{n}).$
@@
Power of KL-regularization
The crucial point of the sharper result is utilizing the strong convexity of the objective $Q$ because of the KL-regularization. Specifically, we take the first-order Taylor expansion of sub-optimality with respect to $\{\Delta(x,a)=R(\widehat{\theta},x,a)-R(\theta^{*},x,a):a\in\mathcal{A}\}$
Now, under Algorithm 1, the coverage condition (Definition 2.6) and with concentration inequalities, if the datasize $m=\Theta(\eta^{2}D^{2})$, we can prove that for $\|R(\widehat{\theta},\cdot,\cdot)-R(\theta^{*},\cdot,\cdot)\|_{\infty}\leq\eta^{-1}$ and $\|R(\widehat{\theta}_{0},\cdot,\cdot)-R(\theta^{*},\cdot,\cdot)\|_{\infty}\leq\eta^{-1}$, which implies the whole-policy coverage condition for all contexts: $\|\pi_{f}^{\eta}(\cdot|\cdot)/\pi_{\widehat{\theta}_{0}}^{\eta}(\cdot|\cdot)\|_{\infty}\leq e^{4}.$ Therefore, by setting $n=\Theta(\eta/\epsilon)$, we obtain that $\pi_{\widehat{\theta}}^{\eta}$ is $O(\epsilon)$ optimal. The conclusion is presented in the following theorem.
@@
3.4 Discussion: Result for Local-Coverage
In this subsection, we consider a more general assumption as described in Definition 2.8.

### v2

@@ 3.1 Problem Setup
$Q(\pi)=\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi(\cdot|x)}\bigg{[}R(\theta_{*},x,a)-\eta^{-1}\log\frac{\pi(a|x)}{\pi_{0}(a|x)}\bigg{]},$ (3.1)
@@
In RLHF with the absolute-rating feedback, we can directly measure the quality of the responses by querying absolute reward value. For instance, in the NVIDIA Helpsteer project (Wang et al., 2023b, 2024c), human labelers are required to provide absolute score in five attributes: helpfulness, correctness, coherence, complexity, and verbosity.
The absolute-rating feedback is directly modeled as the stochastic reward in the contextual bandit setting (Wang et al., 2024a; Xiong et al., 2024b). Under the online RLHF setting, in each round $t$, the learner observes a prompt $x_{t}$ (modeled as the context) and chooses a response $a_{t}$ (modeled as the action). The learner then updates the model (policy) based on the absolute-rating feedback.
Remark 3.2.
@@
Reward function class
@@
Hence, the comparator policy is the solution to the oracle given the true reward function $R(\theta_{*},\cdot,\cdot)$: $\pi^{*}(\cdot|\cdot)\propto\pi_{0}(\cdot|\cdot)\cdot\exp(\eta R(\theta_{*},\cdot,\cdot))$. The goal is to minimize the sub-optimality of our learned policy $\widehat{\pi}$ with respect to $\pi^{*}$: $Q(\pi^{*})-Q(\widehat{\pi})$.
Coverage conditions
It is crucial to assume that our data-collector policy $\pi_{0}$ possesses good coverage, which can ensure that the learned reward function can generalize well to unseen contexts (prompts) and actions (responses), and thus can enable us to approximate the optimal policy.
@@
Given a reference policy $\pi_{0}$, $D^{2}$ is the minimum positive real number satisfying $\forall\ {(x,a)\in\mathcal{X}\times\mathcal{A},\ \pi(a|x)>0}$, we have for any pair of $\theta,\theta^{\prime}\in\Theta$,
$\frac{[R(\theta^{\prime},x,a)-R(\theta,x,a)]^{2}}{\mathbb{E}_{x^{\prime}\sim d_{0},a^{\prime}\sim\pi_{0}(\cdot|x^{\prime})}[(R(\theta^{\prime},x^{\prime},a^{\prime})-R(\theta,x^{\prime},a^{\prime}))^{2}]}\leq D^{2}.$
The coverage coefficient $D$ measures how well the in-sample error induced by distribution $d_{0}\times\pi_{0}$ can characterize the out-of-sample error. This concept is adapted from the F-design for online RL under general function approximation (Agarwal et al., 2024), and follows the coverage coefficient for offline RL (Di et al., 2023; Ye et al., 2024b), and the eluder dimension (Wang et al., 2020; Ye et al., 2023; Agarwal et al., 2023; Zhao et al., 2023a) for online RL. Take the linear model as an example, where the reward function is embedded into a $d$-dimensional vector space: $R(\theta,x,a)=\theta^{\top}\phi(x,a)$ for $\theta\in\mathbb{R}^{d}$. Let the covariance matrix $\Sigma=\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{0}(\cdot|x)}\phi(x,a)\phi(x,a)^{\top}.$ Then, the coverage condition turns into
$\sup_{\theta,\theta^{\prime}\in\Theta}\frac{|(\theta^{\prime}-\theta)^{\top}\phi(x,a)|^{2}}{(\theta^{\prime}-\theta)^{\top}\Sigma(\theta^{\prime}-\theta)}\leq\|\phi(x,a)\|_{\Sigma^{-1}}^{2}\leq D^{2},$
where the first inequality uses the Cauchy-Schwarz inequality. Hence, this quantity measures how much does the reference policy covers all directions of the feature space, and we can show that there exists $\pi_{0}$ such that $D^{2}=O(d)$ through G-optimal design (Zhang, 2023; Lattimore and Szepesvári, 2020).
@@ 3.2 Lower Bound
In this subsection, we provide a lower bound for the KL-regularized contextual bandit problem.
@@
For any $\epsilon\in(0,1/256),\eta>4$, and any algorithm $A$, there exists a KL-regularized contextual bandit problem with reward function class $\mathcal{R}$ and $O(N_{\mathcal{R}}(\epsilon))$ data coverage coefficient (as defined in Definition 3.5) such that $A$ requires at least $\Omega\big{(}\min(\frac{\eta\log N_{\mathcal{R}}(\epsilon)}{\epsilon},\frac{\log N_{\mathcal{R}}(\epsilon)}{\epsilon^{2}})\big{)}$ rounds to achieve a suboptimality gap of $\epsilon$.
@@
The lower bound in Theorem 3.6 indicates that the sample complexity of the KL-regularized contextual bandit problem is $\Omega(\eta\log N_{\mathcal{R}}(\epsilon)/\epsilon)$ when $\epsilon$ is sufficiently small. In our proof, the KL-regularization term shifts the local landscape of the objective function, which prevents us to directly apply the standard bandit analysis, and thus requires a novel analysis to derive the new lower bound. This $\Omega(\eta\log N_{\mathcal{R}}(\epsilon)/\epsilon)$ lower bound suggests that the KL-regularized contextual bandit problem potentially enjoy a lower sample complexity compared to the standard contextual bandit.
@@ 3.3 The Proposed Algorithm
We present our algorithm in Algorithm 1 for the KL-regularized contextual bandit problem, which serves as a theoretical model for online RLHF with absolute-rating feedback. The algorithm consists of two stages:
  • In the second stage, we utilize the trained policy $\pi_{\widehat{\theta}_{0}}^{\eta}$ to sample $n$ contexts (prompts) and actions (responses). With the intermediate policy $\pi_{\widehat{\theta}_{0}}^{\eta}$, we can collect new data $\{(x_{i},a_{i},r_{i})\}_{i=1}^{n}$ which is more aligned with the data distribution induced by the optimal policy $\pi_{*}$. In line 15, the algorithm combines data from both stages $\{(x_{i},a_{i},r_{i})\}_{i=1}^{n}$ and $\{(x_{i}^{0},a_{i}^{0},r_{i}^{0})\}_{i=1}^{m}$ to compute a refined least squares estimate $\widehat{\theta}$ of the reward function, minimizing the sum of squared errors across both datasets. By aggregating the two datasets together, there is an overlap between the data to compute $\widehat{\theta}$ and $\widehat{\theta}_{0}$, so that the output policy $\pi_{\widehat{\theta}}^{\eta}$ is well covered by the intermediate policy $\pi_{\widehat{\theta}_{0}}^{\eta}$.
Algorithm 1 Two-stage Mixed-Policy Sampling (TMPS)
@@
10:
11:for $i$ = $1,\ldots,n$ do
12:     Sample context $x_{i}\sim d_{0}$ and action $a_{i}\sim\pi_{\widehat{\theta}_{0}}^{\eta}(\cdot|x_{i})$.
13:     Observe reward $r_{i}=R(\theta_{*},x_{i},a_{i})+\epsilon_{i}$, where $\epsilon_{i}$ is the random noise.
14:end for
15:Compute the least square estimate of the reward function using $\{(x_{i},a_{i},r_{i})\}_{i=1}^{n}$ together with $D_{0}$:
@@
16:Output $\pi_{\widehat{\theta}}^{\eta}(\cdot|\cdot)\propto\pi_{0}(\cdot|\cdot)\exp\bigl{(}\eta R(\widehat{\theta},\cdot,\cdot)\bigr{)}$.
@@ 3.4 Theoretical Guarantees
Review of previous analysis
The previous analysis (e.g., Xiong et al., 2024a) basically follows the techniques of bandits and neglects the significance of KL-regularization. For simplicity, We use short-hand notation $R(\theta,x,\pi)=\mathbb{E}_{a\sim\pi(\cdot|x)}R(\theta,x,a)$ and denote $\mathrm{KL}(\pi(\cdot|x)\|\pi^{\prime}(\cdot|x))$ by $\mathrm{KL}(\pi\|\pi^{\prime})$ when there is no confusion. We make the estimation on a dataset $\{(x_{i},a_{i},r_{i}):x_{i}\sim d_{0},a_{i}\sim\pi_{0}(\cdot|x_{i})\}_{i=1}^{n}$: $\pi_{\widehat{\theta}}^{\eta}=\mathop{\mathrm{argmax}}_{\pi\in\Pi}\mathbb{E}_{x\sim d_{0}}[R(\widehat{\theta},x,\pi)-\eta^{-1}\mathrm{KL}(\pi\|\pi_{0})]$, and has a small in-sample-error: $\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{o}(\cdot|x)}\big{[}(R(\widehat{\theta},x,a)-R(\theta_{*},x,a))^{2}\big{]}=O(1/n).$ The sub-optimality is decomposed as:
$Q(\pi^{*})-Q(\pi_{\widehat{\theta}}^{\eta})$ $=\mathbb{E}_{x\sim d_{0}}\big{[}R(\theta_{*},x,\pi^{*})-R(\widehat{\theta},x,\pi^{*})\big{]}+\mathbb{E}_{x\sim d_{0}}\big{[}R(\widehat{\theta},x,\pi_{\widehat{\theta}}^{\eta})-R(\theta_{*},x,\pi_{\widehat{\theta}}^{\eta})\big{]}$
$+\mathbb{E}_{x\sim d_{0}}\big{[}R(\widehat{\theta},x,\pi^{*})-\eta^{-1}\mathrm{KL}(\pi^{*}\|\pi_{0})\big{]}-\mathbb{E}_{x\sim d_{0}}\big{[}R(\widehat{\theta},x,\pi_{\widehat{\theta}}^{\eta})-\eta^{-1}\mathrm{KL}(\pi_{\widehat{\theta}}^{\eta}\|\pi_{0})\big{]}$
$\leq\mathbb{E}_{x\sim d_{0}}\big{[}R(\theta_{*},x,\pi^{*})-R(\widehat{\theta},x,\pi^{*})+R(\widehat{\theta},x,\pi_{\widehat{\theta}}^{\eta})-R(\theta_{*},x,\pi_{\widehat{\theta}}^{\eta})\big{]},$
@@
$Q(\pi^{*})-Q(\pi_{\widehat{\theta}}^{\eta})\leq$ $2C_{\mathrm{GL}}\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{0}(\cdot|x)}\big{[}|R(\theta_{*},x,a)-R(\widehat{\theta},x,a)|\big{]}$
$\leq$ $2C_{\mathrm{GL}}\sqrt{\mathbb{E}_{a\sim\pi_{0}(\cdot|x)}\big{[}(R(\theta_{*},x,a)-R(\widehat{\theta},x,a))^{2}\big{]}}=O(C_{\mathrm{GL}}/\sqrt{n}).$
@@
Sharper results and analysis
@@
To illustrate the novel techniques to obtain the sharper bound, we highlight the crucial points in the sequel and defer the detailed proof to Appendix C.2.
Part I: Decomposition of the suboptimality gap
The most challenging part is how to proceed with the suboptimality gap based on the strong convexity of the objective $Q$ with the KL-regularization. Given the closed-form solution of $\pi^{*}(a|x)=\pi_{0}(a|x)\exp(\eta R(\theta,x,a))/Z_{\theta_{*}}^{\eta}(x)$ and $\pi_{\widehat{\theta}}^{\eta}(a|x)=\pi_{0}(a|x)\exp(\eta R(\widehat{\theta},x,a))/Z_{\widehat{\theta}}^{\eta}(x)$, where $Z_{\theta}^{\eta}(x)=\sum_{a\in\mathcal{A}}\pi_{0}(a|x)\cdot\exp(\eta R(\theta,x,a))$ denotes the normalization constant, we can write the suboptimality as
$\mathbb{E}_{\pi^{*}}\bigg{[}R(\theta_{*},x,a)-\frac{1}{\eta}\log\frac{\pi^{*}(a|x)}{\pi_{0}(a|x)}\bigg{]}-\mathbb{E}_{\pi_{\widehat{\theta}}^{\eta}}\bigg{[}R^{*}(x,a)-\frac{1}{\eta}\log\frac{\pi_{\widehat{\theta}}^{\eta}(a|x)}{\pi_{0}(a|x)}\bigg{]}$
$\qquad=\frac{1}{\eta}\mathbb{E}_{\pi^{*}}\Big{[}\log\frac{\pi_{0}(a|x)\exp(\eta R(\theta_{*},x,a))}{\pi^{*}(a|x)}\Big{]}-\frac{1}{\eta}\mathbb{E}_{\pi_{\widehat{\theta}}^{\eta}}\Big{[}\log\frac{\pi_{0}(a|x)\exp(\eta R(\theta_{*},x,a))}{\pi_{\widehat{\theta}}^{\eta}(a|x)}\Big{]}$
$\qquad=-\frac{1}{\eta}\big{(}J(x;\widehat{\theta})-J(x;\theta_{*})\big{)},$
where we define $J(x;\theta)=\log Z_{\theta}^{\eta}(x)-\eta\mathbb{E}_{\pi_{\theta}^{\eta}}[R(\theta,x,a)-R(\theta_{*},x,a)]$, and the last equation is deduced by taking the distribution of $\pi^{*}$ and $\pi_{\widehat{\theta}}^{\eta}$ in the terms.
Thus, the suboptimality is expressed by the gap between $\widehat{\theta}$ and $\theta_{*}$ with respect to the function $J$. By taking the first-order Taylor expansion with respect to $\{\Delta(x,a)=R(\widehat{\theta},x,a)-R(\theta_{*},x,a):a\in\mathcal{A}\}$, we can prove the following lemma.
Lemma 3.9.
For any estimator $\widehat{\theta}\in\Theta$, and the policy $\pi_{\widehat{\theta}}^{\eta}$ satisfying Definiton 3.4, we have
$\leq\eta\mathbb{E}_{x\sim d_{0}}\Big{[}\sum_{a\in\mathcal{A}}\pi_{f}^{\eta}(a|x)\Delta^{2}(x,a)\Big{]},$ (3.2)
The proof of this lemma is provided in Appendix C.2.
Part II: Utilizing the coverage condition
In Algorithm 1, with the coverage condition (Definition 3.5) and the concentration inequalities, if the sample size $m=\Theta(\eta^{2}D^{2})$, we can prove that $\|R(\widehat{\theta},\cdot,\cdot)-R(\theta_{*},\cdot,\cdot)\|_{\infty}\leq\eta^{-1}$ and $\|R(\widehat{\theta}_{0},\cdot,\cdot)-R(\theta_{*},\cdot,\cdot)\|_{\infty}\leq\eta^{-1}$, which implies the whole-policy coverage condition for all contexts: $\|\pi_{f}^{\eta}(\cdot|\cdot)/\pi_{\widehat{\theta}_{0}}^{\eta}(\cdot|\cdot)\|_{\infty}\leq e^{4}.$ Therefore, substituting it back into (3.9) leads to
$Q(\pi^{*})-Q(\pi_{\widehat{\theta}}^{\eta})\lesssim\eta\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a\sim\pi_{\widehat{\theta}_{0}}^{\eta}}\Big{[}\big{(}R(\widehat{\theta},x,a)-R(\theta_{*},x,a)\big{)}^{2}\Big{]}.$
Note that $\widehat{\theta}$ in the RHS is computed using the data sampled from $\pi_{\widehat{\theta}_{0}}^{\eta}$. By setting $n=\Theta(\eta/\epsilon)$, we obtain that $\pi_{\widehat{\theta}}^{\eta}$ is $O(\epsilon)$ optimal.
3.5 Result under Local-Coverage Condition
In this subsection, we consider another coverage conditions appearing in previous work as described in Definition 3.11.
Definition 3.10 (Global-Policy Coverage).
Definition 3.11 (Local KL-ball Coverage, Song et al. 2024).
Remark 3.12 (Relation between Local and Global Coverage Conditions).